OpenStudy (anonymous):

What is the vertex for the graph of y = −3x2 + 18x − 13?

OpenStudy (anonymous):

1. You can plug it into a graphic calculator. 2. The x coordinate for the vertex of y = ax^2 + bx + c is x = -b / (2a). If b = 18 and a = -3, x = -18 / (2•-3) = -18 / -6 = 3 then plug in 3 for x to get the vertex's y coordinate.